Feather Friends, Inc., distributes a high-quality wooden birdhouse that sells for $40 per unit. Variable expenses are $20.00 per unit, and fixed expenses total $180,000 per year. Its operating results for last year were as follows:

Sales

$

1,000,000

Variable expenses

500,000

Contribution margin

500,000

Fixed expenses

180,000

Net operating income

$

320,000

Required:

Answer each question independently based on the original data:

1. What is the product's CM ratio?

2. Use the CM ratio to determine the break-even point in dollar sales.

3. If this year's sales increase by $59,000 and fixed expenses do not change, how much will net operating income increase?

4-a. What is the degree of operating leverage based on last year's sales?

4-b. Assume the president expects this year's sales to increase by 17%. Using the degree of operating leverage from last year, what percentage increase in net operating income will the company realize this year?

5. The sales manager is convinced that a 12% reduction in the selling price, combined with a $60,000 increase in advertising, would increase this year's unit sales by 25%.

a. If the sales manager is right, what would be this year's net operating income if his ideas are implemented?

b. If the sales manager's ideas are implemented, how much will net operating income increase or decrease over last year?

6. The president does not want to change the selling price. Instead, he wants to increase the sales commission by $1.90 per unit. He thinks that this move, combined with some increase in advertising, would increase this year's sales by 25%. How much could the president increase this year's advertising expense and still earn the same $320,000 net operating income as last year?
Answer #1

The answer has been presented in the supporting sheet. For detailed answer refer to the supporting sheet.

Answer Part 1) Contribution Margin Ratio = Contribution margin / sales *100 = 500000/1000000*100 = 50% Thus the correct answe

9 Part 3) 20 Increase in net operating income 21 = increase in sales * contribution margin ratio 2 = 59000*50% +3 = $ 29500 5